OpenStudy (anonymous):

What is the ratio 12 to 63 written as a fraction in lowest terms?

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@563blackghost

OpenStudy (anonymous):

@jabez177

jabez177 (jabez177):

\[\frac{ 12 }{ 63 }\]

jabez177 (jabez177):

What can go into 12 and 63? Basically what times what equals 12 and 63?

OpenStudy (danjs):

put them into prime factors

jabez177 (jabez177):

9 is a good factor so: \[12\div9=?\]

OpenStudy (anonymous):

it equals 1.3

OpenStudy (anonymous):

I think 3

jabez177 (jabez177):

I meant 6 12 divided by 6

jabez177 (jabez177):

\[12\div6\]

OpenStudy (danjs):

factors \[\large\frac{ 2^2*3 }{ 7*3^2 }\] top and bottom have one 3 in common, take those out of both (2^2) / (7*3)

OpenStudy (anonymous):

it equals 2

OpenStudy (danjs):

4 / 21

jabez177 (jabez177):

Hold up. Never mind me. I messed up.

OpenStudy (anonymous):

Its ok Steven But I was Right 12/6=2

jabez177 (jabez177):

\[12\div3=?\]

OpenStudy (anonymous):

4 @jabez177

OpenStudy (danjs):

\[\frac{ 12 }{ 63 }=\frac{ 2*2*3 }{ 7*3*3 }\] both have one 3 in common

jabez177 (jabez177):

Now what you do to the top, you have to do to the bottom. \[63\div3=?\]

OpenStudy (danjs):

leaves \[\frac{ 2*2 }{ 7*3 } = \frac{ 4 }{ 21 }\]

OpenStudy (anonymous):

21 @jabez177

jabez177 (jabez177):

Can \[\frac{ 4 }{ 21 }\] be simplified further?

OpenStudy (anonymous):

no

OpenStudy (danjs):

or you can guess and count up factors to see if they work like that, but if you get some weird numbers , like a 4 digit prime number *13 or something, it gets harder

OpenStudy (anonymous):

Thanks @DanJS and @jabez177 Ya'll both helped a lot and if you guys don't mind can you help me with some more problems but you dont have to

OpenStudy (danjs):

sure i have a few min, post another